OpenStudy (anonymous):

Add and simplify: (x + 2) /x + (x - 2) /x

OpenStudy (shubhamsrg):

what if i ask you 4/7 + 11/7 ?

OpenStudy (anonymous):

what is the end result bro?

OpenStudy (anonymous):

2x over x 2

OpenStudy (anonymous):

qwerty, what do you get when you simplify the middle part? (x+(x-2))

OpenStudy (anonymous):

(x + 2) /x + (x - 2) /x =[(x + 2)+(x - 2)]/x =2x/x =2